OpenStudy (anonymous):

What is the MAD of the data set? {10, 9, 16, 5, 10}

OpenStudy (anonymous):

@israa88

OpenStudy (anonymous):

so can u help me

OpenStudy (anonymous):

do u have options ?

OpenStudy (anonymous):

nope no opitions

OpenStudy (anonymous):

10 + 9 + 16 + 5 + 10 = 50 50 ÷ 5 = 10 10 - 10 = 0 9 - 10 = |-1| = 1 16 - 10 = 6 5 - 10 = |-5| = 5 10 - 10 = 0 0 + 1 + 6 + 5 + 0 = 12 The MAD of this data set is 12

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

no problem sir :)

OpenStudy (black_ninja123):

you forgot to divde the last part

OpenStudy (anonymous):

which part

OpenStudy (black_ninja123):

its 2.4 because 12/5

OpenStudy (anonymous):

12

OpenStudy (black_ninja123):

you have to divde 12 by 5 because there are 5 numbers

OpenStudy (anonymous):

@urbanmorgans did u submit it ?

OpenStudy (anonymous):

no

OpenStudy (black_ninja123):

Step 1 Find the mean. ... Step 2 Find the absolute value of the differences between each value in the data set and the mean. ... Step 3 Find the average of the absolute values of the differences between each value in the data set and the mean.

OpenStudy (anonymous):

so is it 2.4

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

thanks for reminding me @Black_ninja123

OpenStudy (black_ninja123):

Np